# Accelerating Diffusion Models in Offline RL via Reward-Aware Consistency Trajectory Distillation

### v1

@@ 4.1 Offline RL
BC $55.2$ $52.5$ $107.5$ $42.6$ $52.9$ $75.3$ $36.6$ $18.1$ $26.0$ $\ 51.9$
CQL $91.6$ $105.4$ $108.8$ $44.0$ $58.5$ $72.5$ $45.5$ $95.0$ $77.2$ $\ 77.6$
IQL $86.7$ $91.5$ $109.6$ $47.4$ $66.3$ $78.3$ $44.2$ $94.7$ $73.9$ $\ 77.0$
DT $86.8$ $107.6$ $108.1$ $42.6$ $67.6$ $74.0$ $36.6$ $82.7$ $66.6$ $\ 74.7$
TT $95.0$ $110.0$ $101.9$ $46.9$ $61.1$ $79.0$ $41.9$ $91.5$ $82.6$ $\ 78.9$
MOPO $63.3$ $23.7$ $44.6$ $42.3$ $28.0$ $17.8$ $53.1$ $67.5$ $39.0$ $\ 42.1$
MOReL $53.3$ $108.7$ $95.6$ $42.1$ $95.4$ $77.8$ $40.2$ $93.6$ $49.8$ $\ 72.9$
MBOP 105.9 $55.1$ $70.2$ $44.6$ $48.8$ $41.0$ $42.3$ $12.4$ $9.7$ $\ 47.8$
Diffusion QL $\underline{97.2}$$\pm 0.4$ $\underline{112.3}$$\pm 0.8$ $111.2$$\pm 0.9$ $51.5$$\pm 0.3$ $96.6$$\pm 3.4$ $\underline{87.3}$$\pm 0.5$ $48.3$$\pm 0.2$ $\underline{102.0}$$\pm 0.4$ 98.0 ±$\pm 0.5$ 89.389.3\ 89.389.3 $5$
Consistency AC $89.2$$\pm 3.3$ $106.0$$\pm 1.3$ $\underline{111.6}$$\pm 0.7$ 71.9 ±$\pm 0.8$ $\underline{99.7}$$\pm 2.3$ $84.1$$\pm 0.3$ 62.7 ±$\pm 0.6$ $100.4$±$\pm 0.6$ $83.0$$\pm 1.5$ $\underline{89.8}$ $2$
Consistency BC $39.6$$\pm 3.4$ $96.8$$\pm 4.6$ $\underline{111.6}$$\pm 0.7$ $46.2$$\pm 0.4$ $78.3$$\pm 2.6$ $84.1$$\pm 0.3$ $45.4$$\pm 0.7$ $100.4$$\pm 0.6$ $80.8$$\pm 3.4$ 75.975.9\ 75.975.9 $2$
Diffuser $88.9$$\pm 0.3$ $103.3$$\pm 1.3$ $106.9$$\pm 0.2$ $42.8$$\pm 0.3$ $74.3$$\pm 1.4$ $79.6$$\pm 0.55$ $37.7$$\pm 0.5$ $93.6$$\pm 0.4$ $70.6$$\pm 1.6$ 77.577.5\ 77.577.5 $20$
RACTD(Ours) $95.9$$\pm 1.5$ 129.0 ±$\pm 1.3$ 119.6 ±$\pm 1.9$ $\underline{59.3}$$\pm 0.2$ 121.0 ±$\pm 0.5$ 98.9 ±$\pm 1.6$ $\underline{57.9}$$\pm 1.0$ 104.9 ±$\pm 2.1$ $\underline{91.5}$$\pm 2.6$ 97.6 1
@@ Results
As shown in Table 1, RACTD achieves the best or second best performance in almost all task, and outperforms the best baseline in overall average score by a substantial margin. It consistently outperforms the diffusion-based planner Diffuser in all environments, and exceeds the performance of consistency-based actor-critic baseline Consistency AC in all hopper and walker2d tasks.
@@ 4.2 Long Horizon Planning
Next, we showcase the effectiveness of RACTD in complex high-dimensional long-horizon planning and its flexibility to adapt to goal-conditioned tasks. Previously, Diffuser has shown great potential in open-loop long-horizon planning, but requires a significantly larger number of denoising steps compared to closed-loop planning like MuJoCo. We demonstrate that our model can achieve superior performance with a single-step denoising process under the same problem formulation.
@@
We test this ability on D4RL Maze2d (Fu et al., 2020), which is a sparse reward long-horizon planning task where an agent may take hundreds of steps to reach the goal in static environments. Following the setup in Janner et al. (2022), we use a planning horizon $128,256,384$ for U-Maze, Medium and Large respectively. We perform open-loop planning by generating the entire state sequence followed by a reverse dynamics model to infer all the actions from the predicted state sequence. The reward model returns $1$ if the current state reaches the goal and 00 otherwise. The baseline results are reported from Janner et al. (2022) and RACTD results are reported as the mean and standard error of 100 planning seeds.
@@ Results
As shown in Table LABEL:tab:long, both Diffuser and RACTD outperform previous model-free RL algorithms CQL and IQL, and MPPI which usese ground-truth dynamics. Our approach surpasses the diffusion baseline in almost all settings, highlighting its ability to effectively capture complex behavioral patterns and high-dimensional information in the training dataset. Notably, the planning dimension for this task (384 for the Large Maze) is substantially higher than that of MuJoCo tasks (16). As a result, Diffuser requires significantly more denoising steps (256 for the Large Maze) compared to MuJoCo (20 steps). On the contrary, despite the increased task complexity, RACTD still only requires a single denoising step to achieve 11.6×11.6$\times$11.6 × performance boost.
Table 2: (Long-horizon planning) The performance of RACTD, Diffuser, and prior model-free algorithms in the Maze2D environment. The best score is in bold and the second-best is underlined.
@@
MPPI $33.2$ $10.2$ $5.1$ $16.2$
CQL $5.7$ $5$ $12.5$ $7.7$
IQL $47.4$ $34.9$ $58.6$ $47.0$
Diffuser $113.9$$\pm 3.1$ $64$ $1.664$ 121.5 ±$\pm 2.7$ $256$ $4.312$ $123.0$$\pm 6.4$ $256$ $5.568$ $119.5$
@@ 4.3 Inference Time Comparison
Beyond performance improvements, a key contribution of our work is significantly accelerating diffusion-based models for decision-making tasks. The primary computational bottleneck in diffusion models arises from the multiple function evaluations (NFEs) required by the iterative denoising process. By reducing the number of denoising steps to a single NFE, our approach achieves a speed-up roughly proportional to the number of denoising originally required.
@@
To evaluate sampling efficiency, we compare RACTD with different samplers, including DDPM (Ho et al., 2020), DDIM (Song et al., 2020a), and EDM (Karras et al., 2022) using the same network architecture as our teacher model. Additionally, we report the efficiency of Diffuser. Note that since Diffuser employs a different model architecture and generates future state-action pairs, its sampling time may also be influenced by these factors. Table 4 and Table LABEL:tab:long present the wall clock sampling time and NFE for MuJoCo (hopper-medium-replay) and Maze2d. All experiments are conducted on one NVIDIA Tesla V100-SXM2-32GB.
@@ Results
In hopper-medium-replay, RACTD achieved 20×20$\times$20 × reduction in NFEs and a 43×43$\times$43 × speed-up compared to Diffuser. Additionally, our student model requires 80×80$\times$80 × fewer NFEs and samples 142×142$\times$142 × faster than the teacher model. In Maze2d, RACTD significantly accelerates computation compared to Diffuser, achieving approximately 57×57$\times$57 ×, 92×92$\times$92 ×, and 114×114$\times$114 × speed-ups on Umaze, Medium and Large mazes, by reducing NFEs by a factor of $256$ for the Medium and Large mazes. Furthermore, we observe that more denoising steps lead to better model performance. This highlights the advantage of distilling from a slow but high-performing teacher model, which enables better performance compared to training a diffusion model directly with fewer denoising steps.
@@ 5 Ablation Study
We ablate over (1) the impact of the reward objective in both student and teacher model training, (2) the effect of different reward loss weights on model performance and training stability, and (3) the impact of increasing denoising steps with our student model.
@@
RACTD (Ours, Student) 0.015 1 $104.9$
Figure 4: Ablation on reward objective weight.
@@ 5.1 Impact of Reward Objective
Hopper Medium-Replay Unconditioned teacher Reward-Aware teacher
Unconditioned student $50.8$$\pm 0.3$ $96.2$$\pm 0.2$
Reward-Aware student 104.9 ±$\pm 2.1$ $96.0$$\pm 0.3$
NFE Time(s) Score
1 0.0147 $104.9$$\pm 2.1$
2 $0.0241$ 109.8 ±$\pm 0.9$
3 $0.0377$ $108.7$$\pm 0.1$
4 $0.0517$ $107.9$$\pm 1.6$
5.2 Effect of Reward Objective Weights
The reward loss weight is a crucial hyperparameter in our pipeline that impacts both training stability and performance. We plot the reward curve achieved over 200 epochs of student training with reward loss weights ranging from $[0.3,0.5,0.7,1.5]$ on MuJoCo hopper-medium-replay in Figure 4. The mean and standard error are reported across 20 rollouts from intermediate checkpoints.
5.3 Number of Sampling Steps
Since our student model is trained for anytime-to-anytime jumps, it naturally extends to multi-step denoising without additional training. Following the approach in Song et al. (2023), given intermediate denoising timesteps $0<t_{1}<t_{2}<T$, we first denoise from $T$ to 00 as usual. We then add noise again to $t_{1}$ and denoise it back to 00, and repeat this process for $t_{2}$. This iterative refinement can enhance generation quality. We evaluate the student using 2, 3, and 4 denoising steps as reported in Table 5. While multi-step sampling improves performance, we observe that gains do not scale linearly with the number of denoising steps.
@@ 6 Related Work
Diffusion models have emerged as a powerful approach for decision-making tasks in reinforcement learning Janner et al. (2022); Ajay et al. (2022); Wang et al. (2022); Hansen-Estruch et al. (2023); Chi et al. (2023). The integration of diffusion models into RL frameworks typically follows two main paradigms: actor-critic approaches, where diffusion models serve as policy or value networks (Wang et al., 2022; Hansen-Estruch et al., 2023), and policy-based approaches, where diffusion models directly generate action trajectories Janner et al. (2022); Ajay et al. (2022); Chi et al. (2023). While these methods have demonstrated impressive performance on standard RL benchmarks, their practical deployment is hindered by the slow sampling time inherent to vanilla diffusion policies based on DDPM Ho et al. (2020). This limitation poses particular challenges for real-world applications such as robotics, where speed is crucial.
@@
Various approaches have been proposed to accelerate the sampling process in diffusion models. One prominent direction leverages advanced ODE solvers to reduce the number of required denoising steps (Song et al., 2020a; Karras et al., 2022; Lu et al., 2022). Another line of work explores knowledge distillation techniques Luhman and Luhman (2021); Salimans and Ho (2022); Berthelot et al. (2023); Kim et al. (2023); Song et al. (2023), where student models learn to take larger steps along the ODE trajectory. Particularly, consistency models Song et al. (2023) leverage the PFODE to learn direct mappings from noise to sample space. Consistency trajectory models (Kim et al., 2023) enable one-step sampling by learning anytime-to-anytime jumps along the PFODE trajectory, achieving performance that surpasses teacher models.
@@
Consistency models have emerged as a promising policy class for behavior cloning from expert demonstrations in robotics (Lu et al., 2024; Prasad et al., 2024; Wang et al., 2024). In RL, several works have enhanced actor-critic methods by replacing traditional diffusion-based value/policy networks with consistency models, showing faster inference and training speed (Chen et al., 2023; Ding and Jin, 2023; Li et al., 2024). These approaches directly incorporate consistency loss (Song et al., 2023) into the value/policy network training, rather than distilling a separate student model. While Wang et al. made some initial attempts to apply consistency distillation in policy-based RL through classifier-free guidance and reverse dynamics, their approach requires two NFEs and under-performs the state-of-the-art. In contrast, RACTD is a straightforward approach of using a separate reward model and incorporating reward objective during student distillation, achieving superior performance with only one NFE.

### v2

@@ 4.1 Offline RL
Offline model selection
BC $55.2$ $52.5$ $107.5$ $42.6$ $52.9$ $75.3$ $36.6$ $18.1$ $26.0$ 51.9\ 51.9 -
CQL $91.6$ $105.4$ $108.8$ $44.0$ $58.5$ $72.5$ $45.5$ $95.0$ $77.2$ 77.6\ 77.6 -
IQL $86.7$ $91.5$ $109.6$ $47.4$ $66.3$ $78.3$ $44.2$ $94.7$ $73.9$ 77.0\ 77.0 -
DT $86.8$ $107.6$ $108.1$ $42.6$ $67.6$ $74.0$ $36.6$ $82.7$ $66.6$ 74.7\ 74.7 -
$T$ $95.0$ $110.0$ $101.9$ $46.9$ $61.1$ $79.0$ $41.9$ $91.5$ $82.6$ 78.9\ 78.9 -
MOPO $63.3$ $23.7$ $44.6$ $42.3$ $28.0$ $17.8$ $\underline{53.1}$ $67.5$ $39.0$ 42.1\ 42.1 -
MOReL $53.3$ $108.7$ $95.6$ $42.1$ $\underline{95.4}$ $77.8$ $40.2$ $93.6$ $49.8$ 72.9\ 72.9 -
MBOP 105.9 $55.1$ $70.2$ $44.6$ $48.8$ $41.0$ $42.3$ $12.4$ $9.7$ 47.8\ 47.8 -
Diffusion QL $\underline{96.8}$ $\pm 0.3$ $\underline{111.1}$ $\pm 1.3$ $110.1$ $\pm 0.3$ $51.1$ $\pm 0.5$ $90.5$ $\pm 4.6$ $\underline{87.0}$ $\pm 0.9$ $47.8$ $\pm 0.3$ $\underline{101.3}$ $\pm 0.6$ $\underline{95.5}$ $\pm 1.5$ 87.9\ 87.9 55
Consistency AC $84.3$ $\pm 4.1$ $100.4$ $\pm 3.5$ $\underline{110.4}$ $\pm 0.7$ 69.1 $\pm 0.7$ $80.7$ $\pm 10.5$ $83.1$ $\pm 0.3$ 58.7 $\pm 3.9$ $99.7$ $\pm 0.5$ $79.5$ $\pm 3.6$ 85.1¯\ $\underline${85.1} 22
Consistency BC $32.7$ $\pm 1.2$ $90.6$ $\pm 9.3$ $\underline{110.4}$ $\pm 0.7$ $31.0$ $\pm 0.4$ $71.7$ $\pm 8.0$ $83.1$ $\pm 0.3$ $34.4$ $\pm 5.3$ $99.7$ $\pm 0.5$ $73.3$ $\pm 5.7$ 69.7\ 69.7 22
Diffuser $88.9$ $\pm 0.3$ $103.3$ $\pm 1.3$ $106.9$ $\pm 0.2$ $42.8$ $\pm 0.3$ $74.3$ $\pm 1.4$ $79.6$ $\pm 0.55$ $37.7$ $\pm 0.5$ $93.6$ $\pm 0.4$ $70.6$ $\pm 1.6$ 77.5\ 77.5 2020
RACTD(Ours) $88.5$ $\pm 2.1$ 120.2 $\pm 2.6$ 122.3 $\pm 0.3$ $\underline{56.6}$ $\pm 0.6$ 102.4 $\pm 2.4$ 112.8 $\pm 1.3$ $\underline{51.4}$ $\pm 0.2$ 108.3 $\pm 1.1$ 105.2 $\pm 1.8$ 96.4 1
Online model selection
Diffusion QL 97.2 $\pm 0.4$ $\underline{112.3}$ $\pm 0.8$ $111.2$ $\pm 0.9$ $51.5$ $\pm 0.3$ $96.6$ $\pm 3.4$ $\underline{87.3}$ $\pm 0.5$ $48.3$ $\pm 0.2$ $\underline{102.0}$ $\pm 0.4$ $\underline{98.0}$ $\pm 0.5$ 89.3\ 89.3 55
[…]
Consistency BC $39.6$ $\pm 3.4$ $96.8$ $\pm 4.6$ $\underline{111.6}$ $\pm 0.7$ $46.2$ $\pm 0.4$ $78.3$ $\pm 2.6$ $84.1$ $\pm 0.3$ $45.4$ $\pm 0.7$ $100.4$ $\pm 0.6$ $80.8$ $\pm 3.4$ $75.9$ 22
RACTD(Ours) $\underline{95.9}$ $\pm 1.5$ 129.0 $\pm 1.3$ 122.3 $\pm 0.3$ $\underline{59.3}$ $\pm 0.2$ 115.5 $\pm 1.7$ 118.8 $\pm 0.3$ $\underline{57.9}$ $\pm 1.0$ 109.5 $\pm 0.3$ 105.2 $\pm 1.8$ 101.5 1
@@ Results
As shown in Table 1, RACTD achieves the highest average score by a substantial margin and best or second-best performance on $8/9$ tasks in Gym-MuJoCo, with the only exception being a medium-expert dataset where reward guidance is less beneficial. RACTD is also the only planning-based method that achieves single-step sampling compared to consistency model based actor-critic methods (which require double sampling steps) and other diffusion-based planners (which require 20× more sampling steps).
Table 2: (Offline RL: FrankaKitchen) Performance and sampling efficiency (NFE) of RACTD and diffusion based baselines on D4RL FrankaKitchen benchmark. Each cell has two values: the first is offline model selection and the second (in brackets) is online model selection.
[…]
Diffusion QL 60.5 $\pm 6.9$ ($63.7$ $\pm 5.2$) 62.6 $\pm 5.1$ ($66.6$ $\pm 3.3$) $\textbf{61.6}(65.15)$ 55
Consistency AC $38.2$ $\pm 1.8$ ($39.8$ $\pm 1.6$) $45.8$ $\pm 1.5$ ($46.7$ $\pm 0.9$) $42.0(43.3)$ 22
Consistency BC $22.6$ $\pm 3.8$ ($23.8$ $\pm 2.8$) $23.5$ $\pm 1.8$ ($24.3$ $\pm 1.3$) 23.1(24.1)\ 23.1(24.1) 22
Flow Q-learning $47.5$ $\pm 2.8$ ($53.3$ $\pm 3.2$) $46.0$ $\pm 2.2$ ($53.5$ $\pm 1.9$) 46.8(53.4)\ 46.8(53.4) 1
RACTD(Ours) $\underline{59.0}$ $\pm 1.5$ ($63.1$ $\pm 0.1$) $\underline{60.9}$ $\pm 0.6$ ($61.9$ $\pm 0.2$) $\underline{60.0}(62.5)$ 1
As shown in Table 2, RACTD consistently outperforms its counterpart that incorporates an actor-critic framework for consistency models (Consistency AC) or flow-based models (Flow q-learning) across all dataset qualities. On suboptimal partial and mixed datasets where reward guidance is crucial, our method achieves near-equivalent performance to multi-step diffusion approaches with $5\times$ fewer sampling steps. These findings highlight the effectiveness of our reward-aware distillation framework, which combines decoupled training, noise-free reward modeling, and single-step sampling efficiency.
@@ 4.2 Long Horizon Planning
Next, we showcase the expressiveness of the single-step sampling model RACTD in complex high-dimensional long-horizon planning tasks. Previously, Diffuser has shown great potential in open-loop long-horizon planning, but requires a significantly larger number of denoising steps compared to closed-loop planning like MuJoCo. We demonstrate that our model can achieve superior performance with a single-step denoising process under the same problem formulation.
@@
We test this ability on D4RL Maze2d (Fu et al., 2020), which is a sparse reward long-horizon planning task where an agent may take hundreds of steps to reach the goal in static environments. Maze2d medium and large are also shared environments in other benchmarks like OGBench (Park et al., 2024), where it is demonstrated to be even more challenging than higher-dimensional environments like AntMaze. Following the setup in Janner et al. (2022), we use a planning horizon $128,256,384$ for U-Maze, Medium and Large respectively, all conditioned on two observations: the start and goal position. We perform open-loop planning by generating the entire state sequence followed by a reverse dynamics model to infer all the actions from the predicted state sequence. The reward model returns $1$ if the current state reaches the goal and 0 otherwise. The baseline results are reported from Janner et al. (2022) and RACTD results are reported as the mean and standard error of 100 planning seeds.
@@ Results
As shown in Table LABEL:tab:long, both Diffuser and RACTD outperform previous model-free RL algorithms CQL, IQL, Flow Q-learning, and MPPI. Our approach surpasses the Diffuser baseline in all settings, highlighting its ability to effectively capture complex behavioral patterns and high-dimensional information in the training dataset. Notably, the planning dimension for this task (384 for the Large Maze) is substantially higher than that of MuJoCo tasks (16). As a result, Diffuser requires significantly more denoising steps (256 for the Large Maze) compared to MuJoCo (20 steps). On the contrary, despite the increased task complexity, RACTD still only requires a single denoising step to achieve $11.6\times$ performance boost, showcasing its expressiveness. Furthermore, maintaining the speedup inherited from CTD, our RACTD is able to achieve significant performance improvement over CTD, especially on the more challenging long-horizon tasks.
Table 3: (Long-horizon planning) The performance of RACTD, Diffuser, Flow Q-learning and prior model-free algorithms in the Maze2D environment. Flow Q-learning is close loop planning and Diffuser, RACTD are open loop planning.
@@
MPPI $33.2$ - - $10.2$ - - $5.1$ - - $16.2$
CQL $5.7$ - - 55 - - $12.5$ - - $7.7$
IQL $47.4$ - - $34.9$ - - $58.6$ - - $47.0$
Flow Q-learning 106.7 $\pm$ 6.4 1 - 89.7 $\pm$ 1.9 1 - 107.3 $\pm$ 7.1 1 - 101.2
Diffuser $113.9$ $\pm 3.1$ $64$ $1.664$ 121.5 $\pm 2.7$ $256$ $4.312$ $123.0$ $\pm 6.4$ $256$ $5.568$ $119.5$
@@ 4.3 Inference Time Comparison
Beyond performance improvements, thanks to the single-step sampling ability from consistency trajectory distillation, our RACTD significantly accelerates diffusion-based models while maintaining their superior performance for decision-making tasks. The primary computational bottleneck in diffusion models arises from the multiple function evaluations (NFEs) required by the iterative denoising process. By reducing the number of denoising steps to a single NFE, our approach achieves a speed-up roughly proportional to the number of denoising steps originally required.
@@
To evaluate sampling efficiency, we compare RACTD with different samplers, including DDPM (Ho et al., 2020), DDIM (Song et al., 2020a), and EDM (Karras et al., 2022) using the same network architecture as our teacher and student model. Additionally, we report the efficiency of Diffuser. Note that since Diffuser employs a different model architecture and generates future state-action pairs, its sampling time may also be influenced by these factors. Table 5 and Table LABEL:tab:long present the wall clock sampling time and NFE for MuJoCo (hopper-medium-replay) and Maze2d. All experiments are conducted on one NVIDIA Tesla V100-SXM2-32GB.
@@ Results
In hopper-medium-replay, RACTD achieved $20\times$ reduction in NFEs and a $43\times$ speed-up compared to Diffuser. Additionally, our student model requires $80\times$ fewer NFEs and samples $142\times$ faster than the teacher model. In Maze2d, RACTD significantly accelerates computation compared to Diffuser, achieving approximately $57\times$, $92\times$, and $114\times$ speed-ups on Umaze, Medium and Large mazes, by reducing NFEs by a factor of $256$ for the Medium and Large mazes.
@@ 5 Ablation Study
We evaluate the necessity of incorporating reward objectives during student versus teacher training and include extended analyses on reward loss weights and multi-step sampling in Appendix E.3 E.5.
@@
RACTD (Ours, Student) 0.015 1 $109.5$
Hopper Medium-Replay Unconditioned teacher Reward-Aware teacher
Unconditioned student $50.8$ $\pm 0.3$ $96.2$ $\pm 0.2$
Reward-Aware student 109.5 $\pm 0.3$ $96.0$ $\pm 0.3$
@@ 6 Related Work
Diffusion models have emerged as a powerful approach for decision-making tasks in reinforcement learning Janner et al. (2022); Ajay et al. (2022); Wang et al. (2022); Hansen-Estruch et al. (2023); Chi et al. (2023). The integration of diffusion models into RL frameworks typically follows two main paradigms: actor-critic approaches, where diffusion models serve as policy or value networks (Wang et al., 2022; Hansen-Estruch et al., 2023), and policy-based approaches, where diffusion models directly generate action trajectories Janner et al. (2022); Ajay et al. (2022); Chi et al. (2023). While these methods have demonstrated impressive performance on standard RL benchmarks, their practical deployment is hindered by the slow sampling time inherent to vanilla diffusion policies based on DDPM Ho et al. (2020). This limitation poses particular challenges for speed-sensitive real-world applications such as robotics.
@@
Various approaches have been proposed to accelerate the sampling process in diffusion models. One prominent direction leverages advanced ODE solvers to reduce the number of required denoising steps (Song et al., 2020a; Karras et al., 2022; Lu et al., 2022). Another line of work explores knowledge distillation techniques Luhman & Luhman (2021); Salimans & Ho (2022); Berthelot et al. (2023); Kim et al. (2023); Song et al. (2023), where student models learn to take larger steps along the ODE trajectory. Particularly, consistency trajectory models (Kim et al., 2023) enable one-step sampling by learning anytime-to-anytime jumps along the PFODE trajectory, achieving performance that surpasses teacher models.
@@
Consistency models have emerged as a promising policy class for behavior cloning from expert demonstrations in robotics (Lu et al., 2024; Prasad et al., 2024; Wang et al., 2024). In RL, several works have enhanced actor-critic methods by replacing traditional diffusion-based value/policy networks with consistency models, showing faster inference and training speed (Chen et al., 2023; Ding & Jin, 2023; Li et al., 2024). These approaches directly incorporate consistency loss (Song et al., 2023) into the value/policy network training, rather than distilling a separate student model. Park et al. (2025) distilled a flow-based model into a single-step student but still requires concurrent network optimization under the actor critic framework. While Wang et al. made some initial attempts to apply consistency distillation in policy-based RL through classifier-free guidance and reverse dynamics, their approach requires two NFEs and under-performs the state-of-the-art. In contrast, RACTD is a straightforward approach of using a separate reward model and incorporating reward objective during student distillation, achieving superior performance with only one NFE.